OpenStudy (anonymous):

simplify (square root 5) (5 - square root 6)

OpenStudy (anonymous):

\[(\sqrt{5})(5-\sqrt{6})=(\sqrt{5}*5)-(\sqrt{5}*\sqrt{6})\] \[=5\sqrt{5}-\sqrt{30}\]

OpenStudy (anonymous):

=5sqrt(5) - sqrt(30)